Thank you. Find the slope and y - intercept for y = 2/5 x + 4.
Answer by ewatrrr(24785)


Hi,
*Note: the general slope-intercept form for an equation of a line is:
y = mx + b where m is the slope and b the y-intercept.
y = 2/5 x + 4 is in the slope-intercept form
slope = (2/5)
y -intercept is 4. Pt(0,4)on the line